OpenStudy (anonymous):

The surface area of a rectangular prism is found using the formula SA = 2(lw + lh + wh), where l is length, w is width, and h is height. What is the surface area (in square meters) of a rectangular prism with length 5 m, width 12 m, and height 6 m?

OpenStudy (anonymous):

your given: length=5m, width=12m, height=6m substitute this in the formula: \[S.A.=2(5\times 12) + 2(5 \times 6) + 2(12 \times 6)\]